A motorcycle traveling at 60mph overtakes a car traveling at 30mph that had a three-hour head start. How far from the starting point are the two vehicles?
How do I put this into an algebraic equation and solve it? Thanks

Answer by ankor@dixie-net.com(22740)

A motorcycle traveling at 60mph overtakes a car traveling at 30mph that had a three-hour head start.
How far from the starting point are the two vehicles?
How do I put this into an algebraic equation and solve it?
:
Let t = travel time of motorcycle
then
(t+3) = travel time of the car
:
Write a distance equation, dist = speed * time
60t = 30(t+3)
60t = 30t + 90
60t - 30t = 90
t = 90/30
t = 3 hrs for the motorcycle to overtake the car
;
Find this dist: 3 * 60 = 180 mi
Check the dist: 6 * 30 = 180 mi